The sum of two consecutive odd integers is 56. Idenify Both of the numbers?
Answer by Alan3354(21605)

The sum of two consecutive odd integers is 56. Idenify Both of the numbers?
----------------------
Call the smaller integer N.
N + (N+2) = 56
2N + 2 = 56
2N = 54
N = 27
The second is N+2, or 29
27 + 29 does = 56, so it's correct.